OpenStudy (anonymous):

The difference between two integers is 5. If the multiplicative inverse of the smaller is added to twice the multiplicative inverse of the larger the result is 23/66. What are the two integers?

OpenStudy (kropot72):

The multiplicative inverse is the reciprocal. Let the smaller integer = x Let the larger integer = y y - x = 5 Therefore y = 5 - x ...............(1) \[\frac{1}{x}+\frac{2}{y}=\frac{23}{66}\ ............(2)\] To find the value of x, substitute the value of y given in equation (1) into equation (2) and solve.

OpenStudy (anonymous):

I've gotten that far, but I am having trouble after this point.

OpenStudy (kropot72):

Please post what you have so far.

OpenStudy (anonymous):

(1/x-5)+(2/x)=23/66 then I try to find the common denominator which would be 66(x-5)?

OpenStudy (kropot72):

\[\frac{1}{x}+\frac{2}{5-x}=\frac{23}{66}\] The common denominator for the left hand side is x(5-x) \[\frac{5-x+2x}{x(5-x)}=\frac{23}{66}\ ..........(3)\] Now you can cross multiply equation (3) to get rid of the fractions.

OpenStudy (anonymous):

So now I have 330=49x-23x^2 is this correct so far?

OpenStudy (kropot72):

Good work! So now we have a quadratic to solve: \[23x ^{2}-49x+330=0\ ...........(4)\]

OpenStudy (anonymous):

Am I able to just plug this into the quadratic formula?

OpenStudy (kropot72):

That's right :)

OpenStudy (kropot72):

Is the question right. I get a negative disciminant when I try to solve (4).

OpenStudy (kropot72):

*discriminant

OpenStudy (anonymous):

I got: \[x=\frac{ -49\pm i \sqrt{27959} }{ -46 } \]

OpenStudy (anonymous):

I just checked and the answer is actually 11 and 6

OpenStudy (kropot72):

My result is not the same. However we both get imaginary solutions rather than an integer. Please check the question again. Do you have answer choices given or a given answer?

OpenStudy (anonymous):

The given answers are 11 and 6

OpenStudy (kropot72):

Sorry. My bad. Equation (1) should have been y = x + 5 y = x + 5 .................(1) \[\frac{1}{x}+\frac{2}{y}=\frac{23}{66}\ ...........(2)\] Sustituting (1) in (2) gives: \[\frac{1}{x}+\frac{2}{5+x}=\frac{23}{66}\] The quadratic then becomes: \[23x ^{2}-83x-330=0\ ......(3)\] Solving (3) using the quadratic solution formula gives the correct answer fo x (the smaller integer.

OpenStudy (anonymous):

Ok thank you for your help!

OpenStudy (kropot72):

\[\frac{1}{x}+\frac{2}{5+x}=\frac{5+x+2x}{x(5+x)}=\frac{5+3x}{5x+x ^{2}}=\frac{23}{66}\ .........(4)\] Cross multiplying (4) gives: \[330+198x=115x+23x ^{2}\ ..........(5)\] Rearranging (5) gives the quadratic: \[23x ^{2}-83x-330=0\ .........(6)\]

OpenStudy (kropot72):

You're welcome :)